OpenStudy (anonymous):

If AB = 3x + 1 and BC = 6x - 8, find the value of x so that AB = BC.

OpenStudy (anonymous):

\[3x+1=6x-8\] group like terms

OpenStudy (anonymous):

since AB=BC 3x + 1 =6x - 8 i.e. 1+8= 6x-3x 9= 3x x=9/3 x=3